# Visual Prompt Tuning in Null Space for Continual Learning

### v2

@@ 3 Preliminaries
It is worth noting that the rows of the attention map where the prompts serve as queries (i.e., ${\bf Q}_{\bf P}$) do not need to be computed, as formulated in Eq. (4) and illustrated in Figure 1 . The reason is that in VPT-Deep [13], the output prompts of this ViT layer will be omitted, with new trainable prompts introduced in the subsequent layer. If no new prompts are added in the next layer, the output prompts can be just discarded as well.
@@
${{\bf X}}_{t}\mathbf{\Theta}_{t}={{\bf X}}_{t}(\mathbf{\Theta}_{t}+\Delta\mathbf{\Theta})\Rightarrow{{\bf X}}_{t}\Delta\mathbf{\Theta}=\mathbf{0}.$ (8)
@@ 4.1 Analysis of Consistency Conditions
However, the nonlinear operation (i.e., $\mathrm{softmax}$) and the potential higher-order term ${\bf W}_{k}^{\top}{\bf Z}^{\top}{\bf Z}{\bf W}_{v}$ arising from ${\bf K}_{\bf Z}^{\top}{\bf V}_{\bf Z}$ in Eq. (3) complicate the direct resolution of this objective. To address this issue, we propose two sufficient conditions consisting solely of linear operations. Specifically, we split the process of self-attention into two primary stages, i.e., the Affinity described by Eq. (4) and the Aggregation outlined in Eq. (6). We can achieve Eq. (10) by ensuring the consistency of each stage:
@@
Based on Eq. (11--16), we are able to derive the following two equations, respectively:
@@ 4.3 Extension to Multi-Heads
We further extend the consistency conditions Eq. (23) and (24) to multi-head self-attention, a common feature in current transformer-based models. Suppose there are $H$ heads and $d=D/H$ represents the dimension of each token in a head. We use ${\bf Q}_{{\bf X}_{t}.h}\in\mathbb{R}^{N\times d}$, ${\bf W}_{k.h}\in\mathbb{R}^{D\times d}$ and ${\bf S}_{{\bf P}_{t}.h}\in\mathbb{R}^{N\times M}$ to denote the corresponding matrices in Eq. (23) and (24) for the $h$-th head, respectively. The objective is to ensure these conditions are met across all heads, i.e., ${\bf Q}_{{\bf X}_{t}.h}{\bf W}_{k.h}^{\top}\Delta{\bf P}^{\top}=\mathbf{0}$ and ${\bf S}_{{\bf P}_{t}.h}\Delta{\bf P}=\mathbf{0}$, $\forall h\in\{1,2,\cdots,H\}$. Let $\mathbf{\Omega}_{1,t}=[{\bf Q}_{{\bf X}_{t}.1}{\bf W}_{k.1}^{\top};\cdots;{\bf Q}_{{\bf X}_{t}.H}{\bf W}_{k.H}^{\top}]\in\mathbb{R}^{HN\times D}$ and $\mathbf{\Omega}_{2,t}=[{\bf S}_{{\bf P}_{t}.1};\cdots;{\bf S}_{{\bf P}_{t}.H}]\in\mathbb{R}^{HN\times M}$ represent the concatenated matrices from all the heads, respectively. Based on block matrix properties, those two sets of conditions can be formulated as $\mathbf{\Omega}_{1,t}\Delta{\bf P}^{\top}=\mathbf{0}$ and $\mathbf{\Omega}_{2,t}\Delta{\bf P}=\mathbf{0}$. Therefore, for the multi-heads variant, only an additional step of concatenation of the matrices from all heads is required in our algorithm.
@@ 5.2 Main Results
Comparison with Existing Methods: We compare our method with existing methods in Table 2, where the competitors include many recent works. The proposed VPT-NSP2 achieves state-of-the-art accuracies on the four benchmarks, demonstrating the superiority of our method.
@@
More analysis of stability and plasticity, prompt depth, prompt length and memory usage can be seen in the appendix C.
@@ Appendix A Algorithm
The algorithm of our approach is provided in Algorithm 1 . We first initialize the overall uncentered covariance matrices [36] ${\bf C}_{1}$ and ${\bf C}_{2}$, as well as the null-space projection matrices $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. During training, the cross-entropy loss for classification and the loss of prompt distribution $\mathcal{L}_{\mathrm{LN}}$ are jointly utilized for optimization. Subsequently, we get the candidate prompt updates ${\bf P}_{\mathcal{G}}$ computed by the optimizer. Then ${\bf P}_{\mathcal{G}}$ is projected by the null-space projection matrices $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ for updating the prompts. After the convergence, we obtain the matrices ${\bf J}_{1}$ and ${\bf J}_{2}$ to temporarily store ${\bf Q}_{{\bf X}_{t}}{\bf W}_{k}^{\top}$ and ${\bf S}_{{\bf P}_{t}}$ for the data of the current task. Then they are used to update the uncentered covariance matrices ${\bf C}_{1}$ and ${\bf C}_{2}$ by addition. Finally, we update the null-space projection matrices using the uncentered covariance matrices, which will be used in the next task.
Algorithm 2 shows the process of computing a null-space projection matrix. First, an input uncentered covariance matrix ${\bf C}$ is decomposed by SVD, from which we can get the singular values and right singular vectors. Next, we determine the nullity $R$ (i.e., the dimension of null space) of ${\bf C}$ according to the maximum second derivative, which is introduced in Section C.1 . Then we select $R$ right singular vectors corresponding to the $R$ smallest singular values considered close to 0 as the bases of null space. Finally, we compute the normalized projection matrix. In our implementation, the null-space projection matrix is added by an identity matrix with a weight $\eta$ (specifically $\eta_{1}$ for $\mathcal{B}_{1}$ and $\eta_{2}$ for $\mathcal{B}_{2}$). $\eta$ is a hyper-parameter for the trade-off between stability and plasticity, which is also introduced in Section C.1
@@
0:  Uncentered covariance matrix ${\bf C}$, hyper-parameter $\eta\in[0,1]$ for the trade-off between stability and plasticity (mentioned in Section C.1)
@@ Appendix B Experimental Setups and Implementation Details
Models: We validate our approach on the Vision Transformer (ViT) [9] and CLIP [27] models in the experiments, whose backbones are both ViT-Base/16 [9]. The ViT is pre-trained on ImageNet-21k, and we insert 4 prompts into each of the 12 layers for fine-tuning, which is referred to as "VPT" [13]. The classifiers are dependently trained in each task and the orthogonal projection is not applicable to them. All the classifiers from the available tasks are concatenated to make prediction during inference. For the CLIP model pre-trained on the WebImageText, we insert 4 prompts into each of the first 3 layers of the image encoder, while the text encoder is kept frozen. The logit scale that serves as a learnable scalar parameter to scale the cosine similarities between image features and text features is also set to trainable. We observed a serious cross-task confusion among the tasks in the CLIP model. Hence, we follow [43] to utilize the class-wise mean and covariance of previous features extracted before the embedding projection head (i.e., the last linear layer of the image encoder) to refine the projection head, after the prompt tuning stage in each task. Note that this technique is not used for the VPT models.
@@
Implementations Details: For all the datasets and models, the images fed into the models are resized to $224\times 224$ pixels and augmented by AutoAugment [7] during training. For the VPT-based models, we use the Adam optimizer [16] with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and a weight decay of $5\times 10^{-5}$ to train 100 epochs with an initial learning rate of 0.01 and a batch size of 256 on all benchmarks. The learning rate is scaled by a factor of 0.1 at the 50-th and 80-th epoch. Our training losses consist of the cross-entropy loss for classification and the loss $\mathcal{L}_{\mathrm{LN}}$ in Eq. (26) whose coefficient is set to 1. Through cross validation on the validation set, we set the temperatures in the cross-entropy loss to 28, 25, 30 and 30 for the 10-split CIFAR100, 20-split CIFAR100, 10-split ImageNet-R and 10-split DomainNet benchmarks. There are two hyper-parameters $\eta_{1}$ and $\eta_{2}$ used for the trade-off between stability and plasticity in null-space projection as introduced in Section C.1 , and we set both of them to be 0.97, 0.95, 0.94 and 0.95 for the four benchmarks by cross validation.
@@
Appendix C More Model Analysis
C.1 Trade-off between Stability and Plasticity
Adaptive Nullity and Plasticity Enhancement Strategy: Given that the null space of covariance matrix does not always exist in practice, Wang et al. [36] suggest approximating it by selecting the bases whose associated singular values approach zero, where the singular values smaller than a specified multiple (denoted as $\gamma$ in our paper) of the smallest one are selected. However, we experimentally find this strategy and the experience for selecting $\gamma$ are not suitable for prompt tuning in ViTs to determine the nullities $R_{1}$ and $R_{2}$ for the uncentered covariance matrices ${\bf C}_{1}$ and ${\bf C}_{2}$ in Algorithm 1 , which will be introduced afterwards. To solve this problem, we propose an adaptive nullity strategy to determine the nullities in an adaptive manner. Utilizing the characteristic that the curve of descending singular values forms an "L" shape, we divide the curve into two parts by the point where the gradient changes fastest to cover most of the small singular values. It is realized by calculating the maximum second derivative of the points:
@@ C.1 Trade-off between Stability and Plasticity
Motivation of Adaptive Nullity: We take the observation on the 10-split ImageNet-R benchmark as an example to explain our motivation of adaptive nullity. When we use the non-adaptive strategy suggested in [36] and set the multiple $\gamma$ from 10 to 50 according to the experience in CNNs, we find that the nullity is far from sufficient for ${\bf C}_{1}$ yet excessive for ${\bf C}_{2}$. After several attempts, we set the multiples as $3\times 10^{6}$ for ${\bf C}_{1}$ and $3$ for ${\bf C}_{2}$, which are two proper values to cover most of the singular values close to zero. Here we show the singular value curves of ${\bf C}_{1}$ in Figure 6 (a), since its corresponding projection matrix $\mathcal{B}_{1}$ has the dominated effect on the performance. It can be seen that all the singular values are selected in the 11-th layer, which means the whole space is regarded as the null space and thus the forgetting becomes large. While in the 8-th layer, only a small portion of singular values are selected, which is unexpected for the purpose of approximating the groundtruth null space as accurate as possible. The reason for this phenomenon is that the relative scales of singular values differ largely in different layers. Therefore, it is better to use different nullities in different layers according to the singular values in each layer.
Motivated by the "L" shape of the curves, we propose to utilize the point with the maximum second derivative in the curve (as formulated in Eq. (27)) to adaptively capture most of the singular values close to zero. The illustration of our strategy is shown in Figure 6 (b). We can see that in the 8-th layer and 11-th layer, most of the singular values close to 0 can be selected, and the multiple $\gamma$ can be determined automatically even if those $\gamma$s in different layers vary significantly. Although we only illustrate the example on 10-split ImageNet-R, the above observations hold true for other benchmarks.
Figure 6: Singular value curves of ${\bf C}_{1}$ in each layer of VPT-NSP2 at the last task of 10-split ImageNet-R. The orange curves denote those singular values smaller than $\gamma$ multiples of the smallest singular value, whose corresponding right singular vectors will be selected as the bases of null space. Their lengths are equal to the nullities. The subfigure (a) shows the non-adaptive strategy of determining nullities suggested in [36]. The curves with insufficient and excessive nullities are marked in purple and red boxes, respectively. The subfigure (b) shows our adaptive strategy of determining nullities, and the curves of corrected nullities are marked in green boxes.
Effects of Adaptive Nullity and Plasticity Enhancement: We first study how the adaptive nullity affects stability on the four benchmarks by increasing and decreasing the nullities based on the adaptively determined nullity. As shown in Figure 7 , when we use smaller portions of the adaptive nullity which decreases the nullity, or increase the nullity by selecting more higher singular values, the forgetting almost always becomes larger. It indicates the adaptive nullity strategy can select most of the singular values close to zero and thus approximate the groundtruth null space more accurately. As a result, it achieves near-optimal stability which has the near-minimum forgetting. Interestingly, we find that the accuracy increases with the increasing nullity in most cases even if the whole space is taken as the null space (i.e., 200%), which inspires us to design the plasticity enhancement strategy (i.e., Eq. (28)) based on the near-optimal stability for a better trade-off between stability and plasticity.
In Figure 8 , we show the effect of plasticity enhancement on 10-split CIAFR-100 and 10-split ImageNet-R benchmarks where the hyper-parameters $\eta_{1}$ and $\eta_{2}$ are set as the same value and vary from 1 to 0. The accuracy first improves with the increasing plasticity, and then declines due to serious forgetting. Therefore, we can use $\eta_{1}$ and $\eta_{2}$ to enhance the plasticity based on near-optimal stability, so as to make a better trade-off and achieve a higher accuracy. Their values on the four benchmarks are determined by cross validation, which have been introduced in Section B .
We make a comparison with our adaptive nullity and plasticity enhancement strategy with the non-adaptive strategy, whose hyper-parameters $\gamma_{1}$ for ${\bf C}_{1}$ as well as $\gamma_{2}$ for ${\bf C}_{2}$ are chosen elaborately for each benchmark. The result is shown in Table 4 , from which we can see that the adaptive nullity can achieve a higher accuracy with a lower forgetting. When using the plasticity enhancement, despite increased forgetting, the accuracy actually improves rather than declines, indicating that it achieves a better trade-off between stability and plasticity.
Figure 7: Performance curves with regard to varying nullities. 100% means our adaptive nullity, based on which we only use its 1%, 25% and 75% smallest singular values as the nullities. 125$\sim$200indicates that we select extra 25$\sim$100of the rest higher singular values, until all the singular values are selected (i.e., 200%).
Figure 8: Effect of the hyper-parameters $\eta_{1}$ and $\eta_{2}$ which are set as the same value $\bar{\eta}$ varying from 1 to 0. To illustrate the effect of $\bar{\eta}$ as it approaches 1 in detail, we display its performance at the interval of 0.02 from 1 down to 0.9.
Table 4: Effectiveness of adaptive nullity and plasticity enhancement
Strategy 10-S-CIFAR100 20S-CIFAR-100 10S-ImageNet-R 10S-DomainNet
[…]
Non-adaptive nullity 91.41 4.17 89.03 5.22 77.60 5.34 83.20 7.81
Adaptive nullity 91.56 3.09 89.65 4.59 78.45 4.10 83.32 5.92
Adaptive nullity &
plasticity enhancement
91.74 3.28 89.89 4.91 78.88 5.06 83.54 8.54
Figure 9: Result curves with regard to prompt depth. We insert the prompts in the first 1, 3, 6, 9 and all of the 12 layers of the ViT, and each layer contains 4 prompts.
Figure 10: Result curves with regard to prompt length. We use 4, 6, 8, 10 and 12 prompts in each layer, and all the layers are equipped with prompts.
C.2 Effects of Prompt Depth and Length
We conduct experiments on various depths of prompts (i.e., the number of layers to insert prompts) and lengths of prompts (i.e., the number of prompts in each layer) to analyze their effects. The results are shown in Figure 9 and Figure 10 , respectively.
It can be seen from Figure 9 that the prompt depth has a significant effect on both the baseline VPT-seq and our approach VPT-NSP2. With the growth of prompt depth, the accuracy of the baseline first increases and then decreases, while its forgetting increases all the time. It indicates that the gain of accuracy is mainly brought by a proper adaptation to downstream tasks, rather than accumulating the knowledge of new tasks. When more layers are equipped with prompts and thus the baseline model can better adapt to new tasks, a severe forgetting occurs and thus the accuracy drops. In contrast, the VPT-NSP2 can steadily improve the accuracy with the growth of prompt depth, demonstrating it can enhance the ability of continual learning with a more powerful model. We notice that there is a marginal decline of the accuracy on the 10-split DomainNet when the prompt depth increases from 9 to 12, which is a reasonable phenomenon w.r.t prompt depth that also observed in Visual Prompt Tuning [13].
Figure 10 shows that the prompt length has a moderate effect on the performance. With a longer prompt length, the VPT-NSP2 may achieve a better accuracy than the default setting (i.e., 4). The forgetting sometimes increases slightly, which indicates the stability is weaken while the stability is enhanced with the growth of prompt length.
C.3 Memory Analysis
We analyze the extra memory of our approach compared with the sequential fine-tuning baseline. As can be seen in the Algorithm 1, during training in sequentially arrived task $\mathcal{T}_{t}$, we need to store the two uncentered covariance matrices ${\bf C}_{1}\in\mathbb{R}^{D\times D}$ and ${\bf C}_{2}\in\mathbb{R}^{M\times M}$ to hard disk, where $D$ is the dimension of the prompt and $M$ is the number of prompts in a ViT layer. Besides, the null-space projection matrices $\mathcal{B}_{1}\in\mathbb{R}^{D\times D}$ and $\mathcal{B}_{2}\in\mathbb{R}^{M\times M}$, as well as the optimized prompts of the last task ${\bf P}_{t-1}\in\mathbb{R}^{M\times D}$ should be kept on RAM during training. Suppose there are $L$ layers of prompts, and hence the total extra memory is:
$\mathrm{Extra~{}Memory}=2L(D^{2}+M^{2})+LMD$ (29)
Take the VPT-NSP2 as an example, where we insert $L=12$ layers of prompts, each of which contains $M=4$ prompts, and the dimension of each prompt is $D=768$. Therefore, the extra memory is about $14.2$ millions of floating points in this case. The extra memory is constant and does not increase with the number of tasks or classes, which provides an advantage in terms of scalability and efficiency for practical applications.

### v4

@@ 3 Preliminaries
It is worth noting that the rows of the attention map where the prompts serve as queries (i.e., ${\bf Q}_{\bf P}$) do not need to be computed, as formulated in Eq. (4) and illustrated in Figure 1 . The reason is that in VPT-Deep [13], the output prompts of this ViT layer will be replaced with new trainable prompts in the subsequent layer. Omitting ${\bf Q}_{\bf P}$ has no impact on the output image tokens of the ViT layer, as the subsequent Aggregation, LayerNorm and MLP operations are performed independently for each token. If no new prompts are added in the next layer, the output prompts can be just discarded as well.
@@
${{\bf X}}_{t}\mathbf{\Theta}_{t}={{\bf X}}_{t}(\mathbf{\Theta}_{t}+\Delta\mathbf{\Theta}),$ (8)
which can be further simplified as:
${{\bf X}}_{t}\Delta\mathbf{\Theta}=\mathbf{0}.$ (9)
@@ 4.1 Analysis of Consistency Conditions
However, the nonlinear operation (i.e., $\mathrm{softmax}$) and the potential higher-order term ${\bf W}_{k}^{\top}{\bf Z}^{\top}{\bf Z}{\bf W}_{v}$ arising from ${\bf K}_{\bf Z}^{\top}{\bf V}_{\bf Z}$ in Eq. (3) complicate the direct resolution of this objective. Specifically, the non-injection property of the $\mathrm{softmax}$ function causes non-unique solutions. The multiplication between ${\bf K}_{{\bf Z}_{t+1}^{\top}}{\bf V}_{{\bf Z}_{t+1}}$ derives a quadratic term $\mathrm{LN}({\bf P}_{t}+\Delta{\bf P})^{\top}\mathrm{LN}({\bf P}_{t}+\Delta{\bf P})$, which result in difficult optimization for $\Delta{\bf P}$.
To address this issue, we propose two sufficient conditions consisting solely of linear operations. Specifically, we split the process of self-attention into two primary stages, i.e., the Affinity described by Eq. (4) and the Aggregation outlined in Eq. (6). We can achieve Eq. (11) by ensuring the consistency of each stage:
@@
Based on Eq. (12--17), we are able to derive the following two equations, respectively:
@@ 4.3 Extension to Multi-Heads
We further extend the consistency conditions Eq. (24) and (25) to multi-head self-attention, a common feature in current transformer-based models. Suppose there are $H$ heads and $d=D/H$ represents the dimension of each token in a head. We use ${\bf Q}_{{\bf X}_{t}.h}\in\mathbb{R}^{N\times d}$, ${\bf W}_{k.h}\in\mathbb{R}^{D\times d}$ and ${\bf S}_{{\bf P}_{t}.h}\in\mathbb{R}^{N\times M}$ to denote the corresponding matrices in Eq. (24) and (25) for the $h$-th head, respectively. The objective is to ensure these conditions are met across all heads, i.e., ${\bf Q}_{{\bf X}_{t}.h}{\bf W}_{k.h}^{\top}\Delta{\bf P}^{\top}=\mathbf{0}$ and ${\bf S}_{{\bf P}_{t}.h}\Delta{\bf P}=\mathbf{0}$, $\forall h\in\{1,2,\cdots,H\}$. Let $\mathbf{\Omega}_{1,t}=[{\bf Q}_{{\bf X}_{t}.1}{\bf W}_{k.1}^{\top};\cdots;{\bf Q}_{{\bf X}_{t}.H}{\bf W}_{k.H}^{\top}]\in\mathbb{R}^{HN\times D}$ and $\mathbf{\Omega}_{2,t}=[{\bf S}_{{\bf P}_{t}.1};\cdots;{\bf S}_{{\bf P}_{t}.H}]\in\mathbb{R}^{HN\times M}$ represent the concatenated matrices from all the heads, respectively. Based on block matrix properties, those two sets of conditions can be formulated as $\mathbf{\Omega}_{1,t}\Delta{\bf P}^{\top}=\mathbf{0}$ and $\mathbf{\Omega}_{2,t}\Delta{\bf P}=\mathbf{0}$. To sum up, The main difference between single-head and multi-heads is that the parameter update should be orthogonal to the subspace spanned by the concatenation matrices from all heads for multi-heads self-attention. Therefore, for the multi-heads variant, only an additional step of concatenation of the matrices from all heads is required in our algorithm.
@@ 5.2 Main Results
Comparison with Existing Methods: We compare our method with existing methods in Table 2, where the competitors include many recent works. The proposed VPT-NSP2 achieves state-of-the-art performance on the four benchmarks, with surpassing the second best approach by an average of 1.49% in accuracy. The forgetting of our approach is not the lowest, which is reasonable since our approach sacrifices some stability for a better trade-off between stability and plasticity. The outperforming accuracy can demonstrate the superiority of our method.
@@
Table 4: Results for 50 tasks and 100 tasks on CIFAR-100, ImageNet-R and DomainNet datasets. † indicates no plasticity enhancement, and ‡ indicates using the balanced plasticity enhancement where $\bar{\eta}$ is the default value less than 1. Our approach still outperforms other methods in long sequences of tasks.
Method 50S-CIFAR100 50S-ImageNet-R 50S-DomainNet 100S-ImageNet-R 100S-DomainNet
Acc. Forgetting Acc. Forgetting Acc. Forgetting Acc. Forgetting Acc. Forgetting
L2P 76.19 12.06 48.53 12.99 59.45 11.53 38.87 15.26 50.52 17.66
EvoPrompt 76.60 13.86 68.53 10.03 67.68 10.41 61.84 15.84 56.35 21.39
OVOR 65.69 14.28 60.08 5.86 66.27 7.43 40.49 8.12 47.65 8.91
InfLoRA 61.49 13.68 59.02 11.02 69.96 9.51 38.16 15.11 44.32 17.85
EASE 74.47 9.31 68.17 7.76 61.20 10.01 47.55 8.22 33.08 32.14
CPrompt 74.97 7.45 68.47 8.16 67.87 9.36 56.95 10.20 53.73 12.14
VPT-Seq 70.47 29.21 56.38 37.91 58.39 44.79 49.72 45.53 46.39 49.34
VPT-NSP2 81.92 6.56 67.32 6.35 70.13 9.92 59.97 10.07 54.44 11.04
VPT-NSP2 82.98 6.66 69.48 6.51 71.28 11.36 62.23 12.13 57.35 13.82
Long-sequence Continual Learning We experiment on 5 benchmarks under the protocols of 50 tasks and 100 tasks to validate that our approach remains effective even within the context of long-sequence continual learning. The results are presented in Table 4. Despite lacking plasticity enhancement, VPT-NSP2 can outperform existing state-of-the-art approaches and especially surpasses L2P by a large margin. This demonstrates that forgetting is still the predominant factor affecting performance in long sequence of tasks. With the plasticity enhancement, VPT-NSP2 achieves significant increase in accuracy (by 1.1$\sim$2.9%). This demonstrates that our plasticity enhancement is effective in learning new knowledge in long-sequence continual learning.
@@ Appendix A Algorithm
Figure 6: Illustration of our algorithm. The input image tokens with prompts are fed into the ViT layer for forward propagation. During optimization, the gradients of the prompts will be projected into the orthogonal direction to the subspace of the previous task ${\cal T}_{t-1}$. The projected prompt update will be used to update the prompts for anti-forgetting.
An overview and algorithm of our approach are provided in Figure 6 and Algorithm 1 , respectively. We first initialize the overall uncentered covariance matrices [36] ${\bf C}_{1}$ and ${\bf C}_{2}$, as well as the null-space projection matrices $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. During training, the cross-entropy loss for classification and the loss of prompt distribution $\mathcal{L}_{\mathrm{LN}}$ are jointly utilized for optimization. Subsequently, we get the candidate prompt updates ${\bf P}_{\mathcal{G}}$ computed by the optimizer. Then ${\bf P}_{\mathcal{G}}$ is projected by the null-space projection matrices $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ for updating the prompts. After the convergence, we obtain the matrices ${\bf J}_{1}$ and ${\bf J}_{2}$ to temporarily store ${\bf Q}_{{\bf X}_{t}}{\bf W}_{k}^{\top}$ and ${\bf S}_{{\bf P}_{t}}$ for the data of the current task. Then they are used to update the uncentered covariance matrices ${\bf C}_{1}$ and ${\bf C}_{2}$ by addition. Finally, we update the null-space projection matrices using the uncentered covariance matrices, which will be used in the next task.
Algorithm 2 shows the process of computing a null-space projection matrix. First, an input uncentered covariance matrix ${\bf C}$ is decomposed by SVD, from which we can get the singular values and right singular vectors. Next, we determine the nullity $R$ (i.e., the dimension of null space) of ${\bf C}$ according to the maximum second derivative, which is introduced in Section C . Then we select $R$ right singular vectors corresponding to the $R$ smallest singular values considered close to 0 as the bases of null space. Finally, we compute the normalized projection matrix, which provides an upper bound for the scale of the projected gradients and prevents excessive gradient magnitudes. In our implementation, the null-space projection matrix is added by an identity matrix with a weight $\eta$ (specifically $\eta_{1}$ for $\mathcal{B}_{1}$ and $\eta_{2}$ for $\mathcal{B}_{2}$). $\eta$ is a hyper-parameter for the trade-off between stability and plasticity, which is also introduced in Section C
@@
0:  Uncentered covariance matrix ${\bf C}$, hyper-parameter $\eta\in[0,1]$ for the trade-off between stability and plasticity (mentioned in Section C)
@@ Appendix B Experimental Setups and Implementation Details
Models: We validate our approach on the Vision Transformer (ViT) [9] and CLIP [27] models in the experiments, whose backbones are both ViT-Base/16 [9]. The ViT is pre-trained on ImageNet-21k, and we insert 4 prompts into each of the 12 layers for fine-tuning, which is referred to as "VPT" [13]. The classifiers are dependently trained in each task and the orthogonal projection is not applicable to them. All the classifiers from the available tasks are concatenated to make prediction during inference. For the CLIP model pre-trained on the WebImageText, we insert 4 prompts into each of the first 3 layers of the image encoder, while the text encoder is kept frozen. The logit scale that serves as a learnable scalar parameter to scale the cosine similarities between image features and text features is also set to trainable. We observed a serious cross-task confusion among the tasks in the CLIP model. Hence, we follow [43] to utilize the class-wise mean and covariance of previous features extracted before the embedding projection head (i.e., the last linear layer of the image encoder) to refine the projection head, after the prompt tuning stage in each task.
@@
Implementations Details: For all the datasets and models, the images fed into the models are resized to $224\times 224$ pixels and augmented by AutoAugment [7] during training. For the VPT-based models, we use the Adam optimizer [16] with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and a weight decay of $5\times 10^{-5}$ to train 100 epochs with an initial learning rate of 0.01 and a batch size of 256 on all benchmarks. The learning rate is scaled by a factor of 0.1 at the 50-th and 80-th epoch. Our training losses consist of the cross-entropy loss for classification and the loss $\mathcal{L}_{\mathrm{LN}}$ in Eq. (27) whose coefficient is set to 1. Through cross validation on the validation set, we set the temperatures in the cross-entropy loss to 28, 25, 30 and 30 for the 10-split CIFAR100, 20-split CIFAR100, 10-split ImageNet-R and 10-split DomainNet benchmarks. There are two hyper-parameters $\eta_{1}$ and $\eta_{2}$ used for the trade-off between stability and plasticity in null-space projection as introduced in Section C , and we set both of them to be 0.97, 0.95, 0.94 and 0.95 for the four benchmarks by cross validation.
@@
Appendix C Trade-off between Stability and Plasticity
Given that the null space of covariance matrix does not always exist in practice, Wang et al. [36] suggest approximating it by selecting the bases whose associated singular values approach zero, where the singular values smaller than a specified multiple (denoted as $\gamma$ in our paper) of the smallest one are selected. However, we experimentally find this strategy and the experience for selecting $\gamma$ are not suitable for prompt tuning in ViTs to determine the nullities $R_{1}$ and $R_{2}$ for the uncentered covariance matrices ${\bf C}_{1}$ and ${\bf C}_{2}$ in Algorithm 1 , which will be introduced afterwards. To solve this problem, we propose an adaptive nullity strategy to determine the nullities in an adaptive manner. Utilizing the characteristic that the curve of descending singular values forms an "L" shape, we divide the curve into two parts by the point where the gradient changes fastest to cover most of the small singular values. It is realized by calculating the maximum second derivative of the points: